# Leveraging Skills from Unlabeled Prior Data for Efficient Online Exploration

### v2

@@ Appendix D Domain Details
OGBench AntSoccer is a navigation task similar to AntMaze, but with the added complexity where the Ant agent must first travel to the location of a soccer ball, then dribble the soccer ball to the goal location. We benchmark on the two navigate datasets, antsoccer-arena-navigate and antsoccer-medium-navigate.
OGBench Cube and Scene are two manipulation domains with a range of manipulation tasks for each domain. Cube involves using a robot arm to arrange blocks from an initial state to a goal state, which requires pick and place actions to move and/or stack blocks. We benchmark on the two sub-domains, cube-single and cube-double which include one and two blocks, respectively. We also benchmark on the Scene environment with two buttons, a drawer, and a window. The most difficult task requires the agent to perform 8 atomic actions: unlock the drawer and window, open drawer, pick up block, place block in drawer, close drawer, open window, lock drawer and window. We benchmark on all 5 tasks for cube-single, cube-double, and scene, and use the play datasets collected by non-Markovian expert policies with temporally correlated noise.
@@
Visual AntMaze is a benchmark introduced by Park et al. (2023a), where the agent must rely on $64\times 64$ image observations of its surroundings, as well as proprioceptive information including body joint positions and velocities to navigate the maze. In particular, the image is the only way for the agent to locate itself within the maze, so successfully learning to extract location information from the image is necessary for successful navigation. The floor is colored such that any image can uniquely identify a position in the maze. The maze layout is the same as the large layout in the state-based D4RL AntMaze benchmark above and we also use the same additional goals. The reward function and the termination condition are also the same as the state-based benchmark.
@@
Appendix E ICVF Implementation Details and Ablation Experiments for Visual Antmaze
@@
We use the public implementation from the authors of (Ghosh et al., 2023) at https://github.com/dibyaghosh/icvf_release and run the ICVF training for $75,000$ gradient steps to obtain the pre-trained encoder weights, following (Li et al., 2024). Then, we initialize the encoder of the RND network with these weights before online learning. It is worth noting that this is slightly different from the prior work (Li et al., 2024) that initializes both the RND network and the critic network. In Figure 6, we examine the performance of Ours, Online w/ Trajectory Skills, and ExPLORe with and without ICVF. Both of the baselines perform much better with the ICVF initialization, suggesting that ICVF might play an important role in providing more informative exploration signal. Ours, without using ICVF, can already outperform the baselines with ICVF. By both extracting skills from offline data and training with offline data, we are able to learn better from less informative exploration signals. We also observe that initializing the critic with ICVF (as done in the original paper (Li et al., 2024)) helps improve the performance of ExPLORe some, but does not substantially change performance.
Appendix F KL Penalty Ablation
@@
Figure 7: Normalized return on three AntMaze mazes, comparing Ours with a KL regularized alternative (Ours (KL)). We that Ours consistently outperforms Ours (KL) on all three mazes, with initial learning that is at least as fast and significantly improved asymptotic performance. Only Ours is able to meet or surpass the asymptotic performance of ExPLORe on all mazes.
@@
In Figure 7, we compare the performance of Ours with a version of our method that uses a KL-divergence penalty with the state-based prior (as used in a previous skill-based method (Pertsch et al., 2021)), Ours (KL). In Ours, as discussed in Section 4, we borrow the policy parameterization from Haarnoja et al. (2018) and adopt a tanh policy parameterization with entropy regularization on the squashed space. Pertsch et al. (2021) parameterize the higher level policy as a normal distribution and is explicitly constrained to a learned state-dependent prior using a KL-divergence penalty, with a temperature parameter that is auto-tuned to match some target value by using dual gradient descent on the temperature parameter. They do not use entropy regularization. Keeping everything else about our method the same, we instantiate this alternative policy parameterization in Ours (KL). We sweep over possible target KL-divergence values $(5,10,20,50)$ and initial values for the temperature parameter $(100,1,0.1)$ using the performance on antmaze-large, but find that these parameters do not substantially alter performance. As shown in Figure 7, Ours performs at least as well as Ours (KL) in the initial learning phase, and has better asymptotic performance on all three mazes, matching or beating ExPLORe, on all three mazes. It seems likely that not having entropy regularization makes it difficult to appropriately explore online, and that explicitly constraining to the prior may prevent further optimization of the policy. Attempts at combining an entropy bonus and KL-penalty lead to instability and difficulty tuning two separate temperature parameters. Additionally, in the Kitchen domain, the KL objective is unstable, since at some states the prior standard deviation is quite small, leading to numerical instability. In contrast, adopting the tanh policy parameterization from Haarnoja et al. (2018) is simple, performs better, and encounters none of these issues in our experiments.
Appendix G Sensitivity Analysis
@@
We picked one representative task from each domain: AntMaze Large top right goal, Kitchen Mixed, HumanoidMaze Giant Stitch, Single Cube Task 1, Double Cube Task 1, Scene Task 1, and AntSoccer Medium Navigate. Figure 8 shows the performance of our method with different RND coefficients. We see that an RND coefficent of zero is very bad for performance on the locomotion tasks, and that the best nonzero RND coeffcient varies between environments. For all experiments in the paper, we selected the middle value of eight, and kept it the same across all domains. Figure 9 shows the performance of our method with different skill horizon lengths. We see that while a horizon length of 2 attains better final performance on AntMaze Large Top Right at the cost of slower initial exploration, it performs worse than a horizon length of 4 on all other environments. A longer horizon length of 8 seems like it could be slightly better in Kitchen Mixed and Scene, but performs much worse in all other tasks. We found that using a horizon length of 4 generally worked well on all tasks, so we used this length for all trajectory-skill based experiments in this paper.
@@
Figure 8: Sensitivity analysis for the RND coefficient. RND is essential to strong performance on the locomotion tasks (AntMaze, AntSoccer, HumanoidMaze). The best coefficient varies between tasks. We use the middle value of 8 for all experiments in this paper, scaled accordingly based on horizon length. All curves use 4 seeds, and standard error is shown.
@@
Figure 9: Sensitivity analysis for the skill horizon length. A horizon length of 4 generally performs the best across all environments. A horizon length of 2 performs relatively well in AntMaze Large Top Right, but performs poorly in all other environments. A longer horizon length of 8 performs slightly better than a horizon length of 4 in Kitchen Mixed and Scene Task 1, but performs poorly in all other tasks. We used a horizon length of 4 for all experiments in this paper. All curves use 4 seeds, and the standard error is shown.
Appendix H State-based D4RL Results
@@
Maze Layout Goal Location Methods without Pretraining Methods with Pretraining Online + RND ExPLORe Diffusion BC w/ JSRL Online w/ Trajectory Skills Online w/ HILP Skills HILP w/ Offline Data Ours Medium Top Left 71 $\pm$ 5.0 27 $\pm$ 3.2 60 $\pm$ 8.1 21 $\pm$ 4.1 120 $\pm$ 47 27 $\pm$ 6.2 14 $\pm$ 3.1 Top Right 100 $\pm$ 16 29 $\pm$ 2.8 85 $\pm$ 19 76 $\pm$ 26 160 $\pm$ 40 72 $\pm$ 36 22 $\pm$ 3.2 Bottom Right 230 $\pm$ 38 35 $\pm$ 4.9 99 $\pm$ 15 77 $\pm$ 34 300 $\pm$ 0 270 $\pm$ 33 22 $\pm$ 4.4 Center 210 $\pm$ 32 71 $\pm$ 8.0 260 $\pm$ 28 26 $\pm$ 3.4 260 $\pm$ 28 300 $\pm$ 0.0 18 $\pm$ 1.7 Aggregated 150 $\pm$ 14 40 $\pm$ 2.0 130 $\pm$ 10 50 $\pm$ 11 210 $\pm$ 17 170 $\pm$ 13 19 $\pm$ 1.8 Large Top Left 72 $\pm$ 10 33 $\pm$ 2.9 52 $\pm$ 3.3 22 $\pm$ 4.2 300 $\pm$ 0 300 $\pm$ 0.0 21 $\pm$ 2.8 Top Right 220 $\pm$ 20 49 $\pm$ 7.7 220 $\pm$ 28 190 $\pm$ 27 280 $\pm$ 20 110 $\pm$ 36 27 $\pm$ 2.6 Bottom Right 280 $\pm$ 15 34 $\pm$ 1.8 160 $\pm$ 22 140 $\pm$ 22 280 $\pm$ 21 260 $\pm$ 19 21 $\pm$ 1.8 Top Center 220 $\pm$ 28 48 $\pm$ 5.2 120 $\pm$ 8.8 59 $\pm$ 12 240 $\pm$ 23 33 $\pm$ 8.5 39 $\pm$ 6.2 Aggregated 200 $\pm$ 8.9 41 $\pm$ 2.7 140 $\pm$ 13 100 $\pm$ 13 270 $\pm$ 12 180 $\pm$ 13 27 $\pm$ 1.7 Ultra Top Left 76 $\pm$ 7.0 34 $\pm$ 4.9 91 $\pm$ 11 36 $\pm$ 11 39 $\pm$ 21 15 $\pm$ 5.3 17 $\pm$ 3.6 Top Right 300 $\pm$ 0.0 92 $\pm$ 20 290 $\pm$ 7.8 120 $\pm$ 14 260 $\pm$ 19 150 $\pm$ 32 37 $\pm$ 5.5 Bottom Right 300 $\pm$ 0.0 70 $\pm$ 8.0 300 $\pm$ 0.0 130 $\pm$ 16 240 $\pm$ 28 67 $\pm$ 12 34 $\pm$ 6.0 Top Center 230 $\pm$ 35 29 $\pm$ 5.5 230 $\pm$ 29 75 $\pm$ 16 100 $\pm$ 32 17 $\pm$ 1.7 22 $\pm$ 4.4 Aggregated 230 $\pm$ 9.3 56 $\pm$ 5.4 230 $\pm$ 9.1 90 $\pm$ 7.1 160 $\pm$ 14 61 $\pm$ 8.2 27 $\pm$ 2.4 Aggregated 190 $\pm$ 6.9 46 $\pm$ 2.3 160 $\pm$ 6.3 80 $\pm$ 5.9 210 $\pm$ 11 130 $\pm$ 7.4 25 $\pm$ 1.4 Table 3: The number of environment steps (×$\times 10^{3}$) taken before the agent find the goal. Lower is better. The first goal time is considered to be $300\times 10^{3}$ steps if the agent never finds the goal. We see that our method is the most consistent, achieving performance as good as or better than all other methods in each of the 4 goals across 3 different maze layouts. The error quantity indicated is standard error over 8 seeds. The method that has the lowest mean is in bold and all the other methods with values that are not statistically significantly higher are also in bold. We used the $t$-test with $p=0.05$ to determine the statistical significance.
@@
Figure 10: Normalized return on individual AntMaze and Kitchen tasks. Ours achieves the strongest performance on all tasks. Online w/ Trajectory Skills learns much slower on all AntMaze tasks, and is asymptotically worse on two of the three more challenging Kitchen tasks. ExPLORe struggles to learn on Kitchen, and performs worse as maze size increases. None of the other baselines are competitive on any tasks. Each curve is an average over four goals with 8 seeds for AntMaze, and 16 seeds for Kitchen.
H.1 Exploration Efficiency
@@
Figure 13 shows the percentage of the maze that the agent has covered throughout the training. The coverage of skill-based methods that do not use prior data during online learning, Online w/ Trajectory Skills and Online w/ HILP Skills, significantly lags behind baselines that use offline data after 50,000 environment steps. Many methods achieve similar coverage on antmaze-medium, likely because the maze is too small to differentiate the different methods. Ours is able to achieve the highest coverage on the antmaze-ultra, and is only surpassed on antmaze-large by HILP w/ Offline Data, which has high first goal times and slow learning. Thus, the coverage difference can likely be at least partially attributed to HILP w/ Offline Data struggling to find the goal and continuing to explore after finding the goal. All non-skill based methods struggle to get competitive coverage levels on antmaze-large and antmaze-ultra. This suggests both pretraining skills and the ability to leverage prior data online are crucial for efficient exploration, and our method effectively compounds their benefits.
H.2 Full D4RL AntMaze Results
@@
We evaluate the success rate of the our algorithm compared to the same baseline suite as in the main results section for each individual goal and maze layout and report the results in Figure 11. We also include ExPLORe both with and without an online RND bonus. Online RND helps ExPLORe the most for the antmaze-medium bottom-right goal, where there is sparse offline data coverage for a considerable radius around the goal. We hypothesize that with the absence of online RND, the agent is encouraged to only stay close to the offline dataset, making it more difficult to find goals in less well-covered regions. On the flip side, for some other goals with better offline data coverage, like the antmaze-large top-right goal, online RND can make the performance worse. For every goal location, Ours consistently matches or outperforms all other methods throughout the training process.
@@
Figure 11: Success rate by goal location. The addition of online RND in ExPLORe leads to better performance on goals with less offline data coverage, and slightly worse performance on goals well-represented in the dataset. Ours consistently matches are outperforms all other methods on all goals throughout training.
@@
We also evaluate the coverage at every goal location for every method for each maze layout and show the result in Figure 12. The coverage varies from goal location to goal location as some goal locations are harder to reach. Generally, the agent stops exploring once it has learned to reach the goal consistently. Ours consistently has the best initial coverage for 11 out of 12 goals, though sometimes has lower coverage compared to other methods later in training. However, this is likely due in large part to successfully learning how to reach that goal quickly, and thus not exploring further.
@@
Figure 12: Coverage for every goal location on three antmaze environments. There is significant variation between goals, and Ours consistently has the best initial coverage performance on 11 of 12 goals. Flattening coverage compared to other methods can be at least partially attributed to having already found the goal, and sucessfully optimizing reaching that goal, rather than continuing to explore after already finding the goal.
@@
Figure 13: Coverage on three different AntMaze mazes, averaged over runs on four goals. Ours has the best coverage performance on the challenging antmaze-ultra, and is only passed by HILP w/ Offline Data on antmaze-large. Online w/ Traj. Skills and Online with HILP Skills struggle to explore after initial learning, and Online and Diffusion BC + JSRL generally perform poorly at all time steps.
H.3 D4RL Play Dataset

### v4

@@ Appendix E Domain Details
OGBench antsoccer is a navigation task similar to antmaze, but with the added complexity where the Ant agent must first travel to the location of a soccer ball, then dribble the soccer ball to the goal location. We benchmark on the two navigate datasets, antsoccer-arena-navigate and antsoccer-medium-navigate. We only benchmark on task-1.
OGBench cube-* and scene are manipulation domains with a range of manipulation tasks for each domain. Cube domains involve using a robot arm to arrange blocks from an initial state to a goal state, which requires pick and place actions to move and/or stack blocks. We use both the cube-single and cube-double domains which include one and two blocks, respectively. We also benchmark on the Scene environment with two buttons, a drawer, and a window. The most difficult task requires the agent to perform 8 atomic actions: unlock the drawer and window, open drawer, pick up block, place block in drawer, close drawer, open window, lock drawer and window. We benchmark on all 5 tasks for cube-single, cube-double, and scene, and use the play datasets collected by non-Markovian expert policies with temporally correlated noise. Unlike the reward function used in the singletask versions of these tasks introduced after this paper was finished, we use a binary -1/0 reward on the completion of the entire episode to create a maximally sparse task. This is different from the reward in the recently introduced singletask versions of these tasks, where the reward is proportional to the number of subtasks completed (1 subtask for cube-single, 2 for cube-double, and 5 for scene).
@@
visual-antmaze is a benchmark introduced by Park et al. (2023a), where the agent must rely on $64\times 64$ image observations of its surroundings, as well as proprioceptive information including body joint positions and velocities to navigate the maze. In particular, the image is the only way for the agent to locate itself within the maze, so successfully learning to extract location information from the image is necessary for successful navigation. The floor is colored such that any image can uniquely identify a position in the maze. The maze layout is the same as the large layout in the state-based D4RL antmaze benchmark above and we also use the same additional goals. The reward function and the termination condition are also the same as the state-based benchmark.
@@
Appendix F ICVF Implementation Details and Ablation Experiments for visual-antmaze
@@
We use the public implementation from the authors of (Ghosh et al., 2023) and run the ICVF training for $75000$ gradient steps to obtain the pre-trained encoder weights, following (Li et al., 2024). Then, we initialize the encoder of the RND network with these weights before online learning. It is worth noting that this is slightly different from the prior work (Li et al., 2024) that initializes both the RND network and the critic network. In Figure 7, we examine the performance of SUPE, Trajectory Skills, and ExPLORe with and without ICVF. Both of the baselines perform much better with the ICVF initialization, suggesting that ICVF might play an important role in providing more informative exploration signal. SUPE, without using ICVF, can already outperform the baselines with ICVF. By both extracting skills from offline data and training with offline data, we are able to learn better from less informative exploration signals. We also observe that initializing the critic with ICVF (as done in the original paper (Li et al., 2024)) helps improve the performance of ExPLORe some, but does not substantially change performance.
Figure 7: Success rate on the visual-antmaze environment with and without ICVF. SUPE works well without ICVF, almost matching the original performance. However, Trajectory Skills achieves far worse performance without ICVF, which shows that using offline data both for extracting skills and online learning leads to better utilization of noisy exploration bonuses. Initializing the critic with ICVF helps for ExPLORe, but does not substantially change performance.
@@
Figure 8: IQM normalized return on three antmaze mazes, comparing SUPE with a KL regularized alternative (Ours (KL)). We that SUPE consistently outperforms Ours (KL) on all three mazes, with initial learning that is at least as fast and significantly improved asymptotic performance. Only SUPE is able to meet or surpass the asymptotic performance of ExPLORe on all mazes.
Appendix G KL Penalty Ablation
@@
In Figure 8, we compare the performance of SUPE with a version of our method that uses a KL-divergence penalty with the state-based prior (as used in a previous skill-based method (Pertsch et al., 2021)), Ours (KL). In SUPE, as discussed in Section 4, we borrow the policy parameterization from (Haarnoja et al., 2018) and adopt a tanh policy parameterization with entropy regularization on the squashed space. Pertsch et al. (2021) parameterize the higher level policy as a normal distribution and is explicitly constrained to a learned state-dependent prior using a KL-divergence penalty, with a temperature parameter that is auto-tuned to match some target value by using dual gradient descent on the temperature parameter. They do not use entropy regularization. Keeping everything else about our method the same, we instantiate this alternative policy parameterization in Ours (KL). We sweep over possible target KL-divergence values $(5,10,20,50)$ and initial values for the temperature parameter $(100,1,0.1)$ using the performance on antmaze-large, but find that these parameters do not substantially alter performance. As shown in Figure 8, SUPE performs at least as well as Ours (KL) in the initial learning phase, and has better asymptotic performance on all three mazes, matching or beating ExPLORe on all three mazes. It seems likely that not having entropy regularization makes it difficult to appropriately explore online, and that explicitly constraining to the prior may prevent further optimization of the policy. Attempts at combining an entropy bonus and KL-penalty lead to instability and difficulty tuning two separate temperature parameters. Additionally, in the Kitchen domain, the KL objective is unstable, since at some states the prior standard deviation is quite small, leading to numerical instability. In contrast, adopting the tanh policy parameterization from (Haarnoja et al., 2018) is simple, performs better, and encounters none of these issues in our experiments.
Appendix H Sensitivity Analysis
@@
We picked one representative task from each domain: antmaze-large-top-right, kitchen-mixed, humanoidmaze-giant-stitch, cube-single-task1, cube-double-task1, scene-task1, and antsoccer-arena-navigate. Figure 9 shows the performance of our method with different RND coefficients. We see that an RND coefficent of zero leads to poor performance on antmaze and humanoidmaze, and that the best nonzero RND coeffcient varies between environments. For all experiments in the paper, we selected the middle value of eight, and kept it the same across all domains. Figure 10 shows the performance of our method with different skill horizon lengths. We see that while a horizon length of 2 attains better final performance on antmaze-large top right goal at the cost of slower initial exploration, it performs worse than a horizon length of 4 on all other environments. A longer horizon length of 8 seems to perform significantly better in soccer and kitchen, but performs the same or substantially worse in all other tasks. We found that using a horizon length of 4 generally works well on all tasks, so we used this length for all trajectory-skill based experiments in this paper.
@@
Figure 9: Sensitivity analysis of the RND coefficient. RND is essential to strong performance antmaze and humanoidmaze. The best coefficient varies between tasks.
@@
Figure 10: Sensitivity analysis of the skill horizon length. A horizon length of 4 generally performs the best across all environments. A horizon length of 2 performs relatively well in antmaze-large, but performs poorly in all other environments. A longer horizon length of 8 performs better than a horizon length of 4 in soccer and kitchen, but performs much worse in several other tasks.
Appendix I State-based D4RL Results
@@
Figure 11: IQM normalized return on individual antmaze and kitchen tasks. SUPE achieves the strongest performance on all tasks. Trajectory Skills learns much slower on all antmaze tasks. ExPLORe struggles to learn on kitchen, and performs worse as maze size increases. None of the other baselines are competitive on any tasks. Each curve is an average over four goals with 8 seeds for antmaze, and 16 seeds for kitchen.
I.1 Exploration Efficiency
@@
Figure 14 shows the percentage of the maze that the agent has covered throughout the training. The coverage of skill-based methods that do not use prior data during online learning, Trajectory Skills and HILP Skills, significantly lags behind baselines that use offline data after $50000$ environment steps. Many methods achieve similar coverage on antmaze-medium, likely because the maze is too small to differentiate the different methods. SUPE is able to achieve the highest coverage on the antmaze-ultra, and is only surpassed on antmaze-large by SUPE (HILP), which has high first goal times and slow learning. Thus, the coverage difference can likely be at least partially attributed to SUPE (HILP) struggling to find the goal and continuing to explore after finding the goal. All non-skill based methods struggle to get competitive coverage levels on antmaze-large and antmaze-ultra. This suggests both pretraining skills and the ability to leverage prior data online are crucial for efficient exploration, and our method effectively compounds their benefits.
I.2 Full D4RL AntMaze Results
@@
We evaluate the success rate of the our algorithm compared to the same baseline suite as in the main results section for each individual goal and maze layout and report the results in Figure 12. We also include ExPLORe both with and without an online RND bonus. Online RND helps ExPLORe the most for the antmaze-medium bottom-right goal, where there is sparse offline data coverage for a considerable radius around the goal. We hypothesize that with the absence of online RND, the agent is encouraged to only stay close to the offline dataset, making it more difficult to find goals in less well-covered regions. On the flip side, for some other goals with better offline data coverage, like the antmaze-large top-right goal, online RND can make the performance worse. For every goal location, SUPE consistently matches or outperforms all other methods throughout the training process.
@@
Figure 12: IQM normalized return by goal location. The addition of online RND in ExPLORe leads to better performance on goals with less offline data coverage, and slightly worse performance on goals well-represented in the dataset. SUPE consistently matches and outperforms all other methods on all goals throughout training.
@@
We also evaluate the coverage at every goal location for every method for each maze layout and show the result in Figure 13. The coverage varies from goal location to goal location as some goal locations are harder to reach. Generally, the agent stops exploring once it has learned to reach the goal consistently. SUPE consistently has the best initial coverage for 11 out of 12 goals, though sometimes has lower coverage compared to other methods later in training. However, this is likely due in large part to successfully learning how to reach that goal quickly, and thus not exploring further.
@@
Figure 13: IQM coverage for every goal location on three antmaze environments. There is significant variation between goals, and SUPE consistently has the best initial coverage performance on 11 of 12 goals. Flattening coverage compared to other methods can be at least partially attributed to having already found the goal, and sucessfully optimizing reaching that goal, rather than continuing to explore after already finding the goal.
@@
Figure 14: IQM coverage on three different AntMaze mazes, averaged over runs on four goals. SUPE has the best coverage performance on the challenging antmaze-ultra, and is only passed by SUPE (HILP) on antmaze-large. Trajectory Skills and HILP Skills struggle to explore after initial learning, and Online Only and DBC+JSRL generally perform poorly at all time steps.
Maze Layout Goal Location Methods without Pretraining Methods with Pretraining Online Only ExPLORe DBC+JSRL Trajectory Skills HILP Skills SUPE (HILP) SUPE medium top-left 71 $\pm$ 5.0 27 $\pm$ 3.2 60 $\pm$ 8.1 21 $\pm$ 4.1 120 $\pm$ 47 27 $\pm$ 6.2 14 $\pm$ 3.1 top-right 100 $\pm$ 16 29 $\pm$ 2.8 85 $\pm$ 19 76 $\pm$ 26 160 $\pm$ 40 72 $\pm$ 36 22 $\pm$ 3.2 bottom-right 230 $\pm$ 38 35 $\pm$ 4.9 99 $\pm$ 15 77 $\pm$ 34 300 $\pm$ 0 270 $\pm$ 33 22 $\pm$ 4.4 center 210 $\pm$ 32 71 $\pm$ 8.0 260 $\pm$ 28 26 $\pm$ 3.4 260 $\pm$ 28 300 $\pm$ 0.0 18 $\pm$ 1.7 Aggregated 150 $\pm$ 14 40 $\pm$ 2.0 130 $\pm$ 10 50 $\pm$ 11 210 $\pm$ 17 170 $\pm$ 13 19 $\pm$ 1.8 large top-left 72 $\pm$ 10 33 $\pm$ 2.9 52 $\pm$ 3.3 22 $\pm$ 4.2 300 $\pm$ 0 300 $\pm$ 0.0 21 $\pm$ 2.8 top-right 220 $\pm$ 20 49 $\pm$ 7.7 220 $\pm$ 28 190 $\pm$ 27 280 $\pm$ 20 110 $\pm$ 36 27 $\pm$ 2.6 bottom-right 280 $\pm$ 15 34 $\pm$ 1.8 160 $\pm$ 22 140 $\pm$ 22 280 $\pm$ 21 260 $\pm$ 19 21 $\pm$ 1.8 top-center 220 $\pm$ 28 48 $\pm$ 5.2 120 $\pm$ 8.8 59 $\pm$ 12 240 $\pm$ 23 33 $\pm$ 8.5 39 $\pm$ 6.2 Aggregated 200 $\pm$ 8.9 41 $\pm$ 2.7 140 $\pm$ 13 100 $\pm$ 13 270 $\pm$ 12 180 $\pm$ 13 27 $\pm$ 1.7 ultra top-left 76 $\pm$ 7.0 34 $\pm$ 4.9 91 $\pm$ 11 36 $\pm$ 11 39 $\pm$ 21 15 $\pm$ 5.3 17 $\pm$ 3.6 top-right 300 $\pm$ 0.0 92 $\pm$ 20 290 $\pm$ 7.8 120 $\pm$ 14 260 $\pm$ 19 150 $\pm$ 32 37 $\pm$ 5.5 bottom-right 300 $\pm$ 0.0 70 $\pm$ 8.0 300 $\pm$ 0.0 130 $\pm$ 16 240 $\pm$ 28 67 $\pm$ 12 34 $\pm$ 6.0 top-center 230 $\pm$ 35 29 $\pm$ 5.5 230 $\pm$ 29 75 $\pm$ 16 100 $\pm$ 32 17 $\pm$ 1.7 22 $\pm$ 4.4 Aggregated 230 $\pm$ 9.3 56 $\pm$ 5.4 230 $\pm$ 9.1 90 $\pm$ 7.1 160 $\pm$ 14 61 $\pm$ 8.2 27 $\pm$ 2.4 Aggregated 190 $\pm$ 6.9 46 $\pm$ 2.3 160 $\pm$ 6.3 80 $\pm$ 5.9 210 $\pm$ 11 130 $\pm$ 7.4 25 $\pm$ 1.4 Table 3: The number of environment steps (×$\times 10^{3}$) taken before the agent find the goal. Lower is better. The first goal time is considered to be $300\times 10^{3}$ steps if the agent never finds the goal. We see that our method is the most consistent, achieving performance as good as or better than all other methods in each of the 4 goals across 3 different maze layouts. The error quantity indicated is standard error over 8 seeds. The method that has the lowest mean is in bold and all the other methods with values that are not statistically significantly higher are also in bold. We use the one-sided $t$-test ($p=0.05$) and bold all the methods that are not statistical significantly worse than the best method (the method that achieves the best mean performance).
I.3 D4RL Play Dataset